Let v1= (1,0,-1), v2=(1,1,1) , v3= (2,2,3). Find the dual basis for {v1,v2,v3).
1
Expert's answer
2021-07-19T17:17:30-0400

By filtering the given vectors into a matrix we have



B = "\begin{bmatrix}
 1 & 1 &2\\
 0 & 1&2\\
-1&1&3
\end{bmatrix}"


B-1 = "\begin{bmatrix}
 1 & -1 &0\\
 -2 & 5&-2\\
1&-2&1\\
\end{bmatrix}"



Now the product B-1 B = I


Hence we see that the the rows of B-1 multiplied by the columns of B satisfy the condition for duality.



Hence, defining



"\gamma_1" = [1, -2, 1]


"\gamma_2" = [-1, 5, -2]


"\gamma_3" = [0, -2, 1]


form the dual basis for {v1, v2, v3} .